OpenStudy (anonymous):

What is the value of time, t, for the given values of distance, d, and rate, r? d = 198 km, r = 36 km/h A.7128 km B.7128 h C.5.5 km D.5.5 h

jimthompson5910 (jim_thompson5910):

distance = rate * time d = r*t d/r = t t = d/r I'll let you finish

OpenStudy (anonymous):

Thanks

jimthompson5910 (jim_thompson5910):

np